# Rel-A.I.: An Interaction-Centered Approach To Measuring Human-LM Reliance

### v1

@@ Abstract
[…]
@@
Kaitlyn Zhou1,4 Jena D. Hwang4 Xiang Ren2,4 Nouha Dziri4 Dan Jurafsky1 Maarten Sap3,4 1Stanford University, 2University of Southern California, 3Carnegie Mellon University, 4Allen Institute for AI katezhou@stanford.edu
@@ 1 Introduction
Figure 1: An interaction-centered approach to measuring human reliance on LM generations with epistemic markers. In situ dimensions of these engagements include: long-term interaction, perceived warmth, and subject matter.
With the rise in popularity of language models (LMs) comes new features and dimensions that characterize human-LM interactions. Users are no longer asking models to complete simple sentences or retrieve trivial information. Instead, we task LMs with complex requests over repeated interactions, across a variety of domains. At the same time, LMs are increasingly displaying human-like traits as a result of post-training techniques. These enhanced capabilities enable richer and more engaging interactions, but LM generations remain susceptible to errors and hallucinations. As the interactions between humans and LMs evolve, there is a need to understand how these new contextual variables, which we call the interactional context, impact human reliance on LMs. Inspired by Suchman’s (2007) work, which imagines the reconfiguration of human-computer interactions and the evaluation paradigm proposed by Lee et al. (2022), we are interested in the reconfiguration of human-LM interactions and the evaluations needed to ensure safe and helpful interactions.
Prior work Kadavath et al. (2022); Tian et al. (2023); Liu et al. (2023a); Xiong et al. (2024); Tanneru et al. (2023); Mielke et al. (2022); Lin et al. (2022); Stengel-Eskin et al. (2024) at the intersection of calibration and human-AI reliance has focused on the verbalized confidence of the system (e.g., “Confidence: 90%” or “I’m certain the answer is…”) as the key factor that contributes to human-LM reliance. In machine learning calibration, as long as each response’s internal probability corresponds with its probability of being correct, then the system is considered calibrated, e.g., expected calibration error Naeini et al. (2015). Similarly, in linguistic calibration Mielke et al. (2022), if the confidence of the verbalized generation matches the probability of the response being correct, then the model is linguistically calibrated.
We contend that human reliance depends on a myriad of factors, with expressed confidence being only one contributing factor. In our work, instead of asking if the model is calibrated or linguistically calibrated, we instead ask if the human-LM interaction is calibrated.
[…]
In our work, we use this methodology to evaluate how prior interactions, the persona of the LM, and the topics discussed might influence human decisions on reliance. First, how do prior interactions with LMs influence reliance behaviors? For example, if a model is typically uncertain (as verbalized) but occasionally expresses moderate confidence, would those statements of moderate confidence be relied on more frequently than if they appeared among a highly confident model (RQ1)? Next, motivated by the rise of humanlike attributes in language models Abercrombie et al. (2021, 2023a), we examine how the warmth of a model, a product of anthropomorphism of LMs Kim et al. (2019) influences human reliance. As public models are fine-tuned to include personas and friendliness, how do these features of anthropomorphism impact human reliance (RQ2)? Lastly, as these language models are being deployed across a variety of domains Zhao et al. (2024); Wang et al. (2024), how might human reliance vary across subject matter (RQ3)?
Our findings illustrate that reliance on LMs is not only contingent on the verbalized confidence but that the interactional context must also be taken into consideration. Prior interactions, anthropomorphism, and subject matter all significantly impact human reliance, with up to 20% change in reliance on the exact same expression as interaction contexts change. For example, statements of moderate confidence appearing in a highly certain model (as verbalized) will be relied on significantly less than the same statement appearing from a typically not confident model. Warm statements like, ”I’m happy to help! I believe the answer is B”, are relied on significantly more often than the same expression without warmth, ”I believe the answer is B”. Generations in computationally heavy domains like math are relied on more often than non-computation subjects like law and philosophy. Our findings consistently show that the exact same expressions of (un)certainty lead to differences in human behavior depending on the context of the interaction.
@@ 2 Background
Machine learning as focused largely on improving model calibration Jiang et al. (2021); Desai and Durrett (2020); Jagannatha and Yu (2020); Kamath et al. (2020); Kong et al. (2020) — aligning the model probabilities with accuracy. Other recent work has examined how pretraining Hendrycks et al. (2019) and scaling Srivastava et al. (2022); Chen et al. (2023b) impact LM calibration. Works in linguistic calibration focused on fine-tuning LMs to be calibrated between what is internally represented and what is verbalized either numerically Kadavath et al. (2022); Tian et al. (2023); Liu et al. (2023a); Xiong et al. (2024); Tanneru et al. (2023), ordinally Mielke et al. (2022); Lin et al. (2022), or verbally Stengel-Eskin et al. (2024). Recent work also explored calibration based on a model’s internal representational states Hofweber et al. (2024).
@@
We ground our work in LMs’ expressions of (un)certainty (e.g., ”I’m certain the answer is…”). These linguistic cues, or epistemic markers, mark speaker stance and commitment and contribute to human sense-making and decision-making Budescu et al. (1988); Windschitl and Wells (1996); Druzdzel (1989). In this work, we are interested in how these expressions of (un)certainty influence humans’ decisions on whether or not to rely on an LM-generated response. [^1] Plain statements lacking epistemic marking (”The answer is…”) are less informative to our study as humans often rely on these responses exclusively and rarely override them Zhou et al. (2024).
@@ 3 Interaction-Centric Approach to Measuring Human-LM Reliance
RQ1 Prior Interactions $S_{1A_{conf}}$ Highly Confident Model 10 20 0
$S_{1B_{unconf}}$ Unconfident Model 0 20 10
RQ2 Degree of Warmth $S_{2A}$ Unnamed Agent (Agent37821) 5 20 5
$S_{2B}$ Named Agent (Gordon) 5 20 5
$S_{2C}$ Named Agent (Marvin) + Warmth 5 20 5
RQ3 Domain $S_{3A}$ Professional Accounting 4 10 4
$S_{3B}$ Professional Psychology 4 10 4
$S_{3C}$ Professional Law 4 10 4
$S_{3D}$ Clinical Knowledge 4 10 4
$S_{3E}$ Trivia Questions 4 10 4
Table 1: Overview of the experimental set-ups where systems vary in prior interactions confidence, the degree of warmth, and subject domain. Each experiment includes 60 or 90 interactions and the number of strengtheners, weakened strengtheners, and weakeners included per system are listed above.
@@ 3.1 Desiderata Of In Situ Evaluation
Our work introduces Rel-A.I., an approach to measure inter-system reliance, with the explicit interest in understanding how interactional context impacts human reliance decisions. We name three key desiderata needed to conduct evaluations of human reliance on LM-generations centered on in situ, system-level, and robust evaluations.
@@
In order to understand how interactional context affects reliance, evaluations must occur in a situated environment. Our evaluations use self-incentivized tasks where users gain points by correctly answering challenging trivia questions, deciding whether to rely on an AI agent’s response for help. [^2] Adopted from Zhou et al. (2024) and Bansal et al. (2019) LMs’ predicted answers are never shown, forcing users to rely on the epistemic markers (e.g., ”I’m certain it’s…”) rather than their own knowledge to make a decision.
@@
A user’s interactional context encompasses their entire interaction with a system and these interactions are rarely isolated to a single event. Humans are known to form mental models of the characteristics (e.g. warmth and competence) and abilities of a system over repeated engagements Norman (1988); Bickmore and Picard (2005). Thus, rather than measuring LM reliance based on multiple interactions with a single system and observing changes over time, we instead have users engage with multiple different systems in one sitting. The systems vary in presentations as required by the experimental setting (i.e., contextual-interaction cues such as warmth of expression or difference in domain knowledge) and the contrast between them enables users to develop separate mental models, allowing us to understand reliance and perception at the system-level.
@@
Humans perceptions of expressions of (un)certainty are internally consistent Druzdzel (1989) but can vary greatly across subjects Budescu et al. (1988); Chesley (1986); Doupnik and Richter (2003). The variance between subjects is a confounding variable that can be expensive to solve (i.e., recruit thousands of participants). However, in our approach, by having the same participant rely on two different systems, we are able to robustly measure intra-subject reliance and use that as a relative score of how contextual interactions influence reliance.
@@
Rel-A.I. consists of three core components, a self-incentivized task, epistemic markers from publicly deployed LMs, and meta-level debrief questions.
@@
Lastly, to meet the desiderata of system-level evaluations, a single participant will partake in multiple rounds of game play with different agents, each with slightly different interactional contexts. The self-incentivized task can be repeated any number of times, enabling researchers the flexibility to test a variety of contexts.
@@
At the center of this methodology is the integration of epistemic markers elicited from publicly deployed LMs into the task. We display to users what language models might actually generate in real-world scenarios and measure how often humans will choose to rely on these generations. In our evaluations, the agent always begins their response with an expression of (un)certainty which cues the participants to the LM’s degree of confidence. We use expressions that were elicited by Zhou et al. (2024) from nine publicly deployed models which were classified into three categories: weakeners — expressions of uncertainty (e.g., ‘‘I think it might be …”, “It could be…”), strengtheners—expressions of certainty (e.g., “I’m certain it’s…”, “Undoubtedly it’s…”), and weakened strengtheners—expressions of moderate certainty (e.g., ”I believe it’s…”, ”Likely it’s…”). See Table 7 and Appendix A.2 for details. [^5] The evaluation of a specific model would be feasible by using the most frequently generated expressions of uncertainty from the model of choice when using the Rel-A.I. approach.
@@ 3.3 Experimental Set Up
With the Rel-A.I. approach, we conduct three experiments, perturbing different interaction variables for each. The first changes the average confidence the models (RQ1), the second varies the degree of warmth of a model (RQ2), and the last diversifies the domain of interaction (RQ3).
@@
Each experiment includes 60 or 90 interactions with different models and variations per system are listed in Table 1. Systems and templates are both presented in random order to the human annotators. We launch the task using Prolific and Qualtrics and recruited 50 new participants for each of our three experimental settings. Our task is compliant with internal review board protocols (Appendix A.1). See A.5 for task screenshots and consent form.
4 Influence of Prior Interactions (RQ1)

### v2

@@ Abstract
The ability to communicate uncertainty, risk, and limitation is crucial for the safety of large language models. However, current evaluations of these abilities rely on simple calibration, asking whether the language generated by the model matches appropriate probabilities. Instead, evaluation of this aspect of LLM communication should focus on the behaviors of their human interlocutors: how much do they rely on what the LLM says? Here we introduce an interaction-centered evaluation framework called Rel-A.I. (pronounced “rely”) that measures whether humans rely on LLM generations. We use this framework to study how reliance is affected by contextual features of the interaction (e.g, the knowledge domain that is being discussed), or the use of greetings communicating warmth or competence (e.g., “I’m happy to help!”). We find that contextual characteristics significantly affect human reliance behavior. For example, people rely 10% more on LMs when responding to questions involving calculations and rely 30% more on LMs that are perceived as more competent. Our results show that calibration and language quality alone are insufficient in evaluating the risks of human-LM interactions, and illustrate the need to consider features of the interactional context.
@@
Kaitlyn Zhou1,4 Jena D. Hwang4 Xiang Ren2 Nouha Dziri4 Dan Jurafsky1 Maarten Sap3,4 1Stanford University, 2University of Southern California, 3Carnegie Mellon University, 4Allen Institute for AI katezhou@stanford.edu
Figure 1: We introduce Rel-A.I., an interaction-centered approach to evaluating LLM risks based on human reliance behaviors. We study the effects of interactional contexts on reliance and find that reliance isn’t solely contingent on the quality of model answer, and that contextual cues such as warmth heavily weigh on human perception of model competence and reliability.
@@ 1 Introduction
With the rise in capabilities of large language models (LLMs), we see an explosion of use cases and tasks for human-AI interactions. [^1] We use the term LLMs except in hyphenated cases where -LM is used to refer to LLMs. The nature of human-LM interactions is also more complex and nuanced than ever before; with users tasking LLMs with complex requests, over repeated interactions, across domains. With new contexts for human-LM interactions and the known overconfidence of language models, a key concern emerges: how might these new contexts introduce risks to the human over-reliance of LLMs?
Prior work has tried to accurately communicate model uncertainties via linguistic Mielke et al. (2022) or numerical calibration Kadavath et al. (2022); Tian et al. (2023); Liu et al. (2023); Tanneru et al. (2023); Lin et al. (2022); Stengel-Eskin et al. (2024); Chaudhry et al. (2024). In both cases, if the confidence of the verbalized generation matches the probability of the response being correct, the model is considered verbally calibrated.
However, our work contends verbal calibration is insufficient to properly measure the safety of human-LM interactions as calibrated generations might still result in unexpected and risky behaviors. Even as models provide risks and hesitations—will humans react with the correct behavior? Instead, we must consider the human behaviors triggered as a result of the generation and evaluate if the generated language triggers the appropriate downstream behaviors.
@@
To address this gap, we introduce an evaluation framework that measures in-situ human reliance on LM-generated responses with epistemic markers called Rel-A.I. (pronounced “rely”) The framework allows us to measure risks of human over-reliance and understand how interactional contexts might influence human decisions. For example, motivated by the rise of humanlike attributes in language models Abercrombie et al. (2021, 2023a), how might the generation of greetings or other social cues influence human reliance (RQ1)? Next, if a model is typically uncertain (as verbalized) but occasionally expresses moderate confidence, would those statements of moderate confidence be relied on more frequently than if they appeared among a highly confident model (RQ2)? Lastly, as these language models are being deployed across a variety of domains Zhao et al. (2024); Wang et al. (2024), how might human reliance vary across subject matter (RQ3) even if models express the same degree of confidence?
Across our experiments, we discover that varying contexts of human-LM interactions can significantly impact human reliance behaviors. Prior interactions, subject matter, and greetings of the model all significantly impact human reliance. For example, statements of moderate confidence appearing in a highly certain model (as verbalized) will be relied on significantly less than the same statement appearing from a typically not confident model. Generations in computational domains like math are relied on more often than non-computational subjects like law and philosophy. Greetings like, ”I’m happy to help! I believe the answer is B” will trigger significantly higher reliance than the same expression without the greeting. Our findings consistently show that the exact same expressions of (un)certainty lead to differences in human behavior depending on the context of the interaction.
At a broader level, our findings demonstrate that models that appear identical in verbal calibration can trigger drastically human behaviors of reliance. Canonical metrics focus on verbal calibration can miss key factors that influence the safety of human-LM interactions. By shifting our measurements toward human reliance, we not only get closer measurements of real-world harms but also evaluate the risks introduced by novel interactional contexts. We urge the research and industry communities to use this approach to evaluate human-LM reliance on their unique language technologies prior to public deployment.
@@ 2 Background
Machine learning as focused largely on improving model calibration Jiang et al. (2021); Desai and Durrett (2020); Jagannatha and Yu (2020); Kamath et al. (2020); Kong et al. (2020)—aligning the model probabilities with accuracy. Other recent work has examined how pretraining Hendrycks et al. (2019) and scaling Srivastava et al. (2022); Chen et al. (2023b) impact LM calibration. Works in linguistic calibration focused on fine-tuning LLMs to be calibrated between what is internally represented and what is verbalized either numerically Kadavath et al. (2022); Tian et al. (2023); Liu et al. (2023); Xiong et al. (2024); Tanneru et al. (2023), ordinally Mielke et al. (2022); Lin et al. (2022), or verbally Stengel-Eskin et al. (2024). Recent work also explored calibration based on a model’s internal representational states Hofweber et al. (2024).
@@
We ground our work in LLMs’ expressions of (un)certainty (e.g., ”I’m certain the answer is…”). These linguistic cues, or epistemic markers, mark speaker stance and commitment and contribute to human sense-making and decision-making Budescu et al. (1988); Windschitl and Wells (1996); Druzdzel (1989). In this work, we are interested in how these expressions of (un)certainty influence humans’ decisions on whether or not to rely on an LM-generated response. [^2] Plain statements lacking epistemic marking (”The answer is…”) are less informative to our study as humans often rely on these responses exclusively and rarely override them Zhou et al. (2024).
@@ 3 Interaction-Centric Approach to Measuring Human-LM Reliance
RQ1 Presentation $A_{control}$ Control Agent 5 20 5
$A_{greet}$ Agent w/ Greeting 5 20 5
RQ2 Prior Interactions $B_{conf}$ Highly Confident Model 10 20 0
$B_{unconf}$ Unconfident Model 0 20 10
RQ3 Domain $C_{account}$ Professional Accounting 4 10 4
$C_{psych}$ Professional Psychology 4 10 4
$C_{law}$ Professional Law 4 10 4
$C_{clinic}$ Clinical Knowledge 4 10 4
$C_{trivia}$ Trivia Questions 4 10 4
Table 1: Overview of the experimental set-ups where systems vary in presentation, prior interactions, and subject domain. Each experiment includes 60 or 90 interactions and the number of strengtheners, weakened strengtheners, and weakeners included per system are listed above.
@@ 3.1 Desiderata Of In Situ Evaluation
Our work introduces Rel-A.I., an evaluation approach that measures human reliance behaviors across various interactional contexts. Below we describe three key features of Rel-A.I. that enable the robust measurement of human reliance on LLM generations, guided by HCI and psycho-linguistics literature.
@@
The gamification of crowdsourced tasks was first introduced by Von Ahn (2006) and is considered a form of intrinsic incentive. Situating users in a game-like scenario not only more closely mimics potential real-world encouragement but has also been shown to encourage high-quality participation Eickhoff et al. (2012); Hossfeld et al. (2014); Krause and Kizilcec (2015); Law et al. (2016). In our task, we situate users in a game-like scenario where users gain points by correctly answering challenging trivia questions, deciding whether to rely on an AI agent’s response for help rather than their own knowledge to make a decision. As we perturb the interactional contexts (i.e., environment of the game), we’re able to measure how each distinct contextual features impact human reliance. [^3] Adopted from Zhou et al. (2024) and Bansal et al. (2019) LLMs’ predicted answers are never shown, forcing users to rely on the epistemic markers (e.g., ”I’m certain it’s…”)
@@
Humans are known to form mental models of the characteristics (e.g. warmth and competence) and abilities of a system over repeated engagements Norman (1988); Bickmore and Picard (2005). Thus, rather than measuring LLM reliance based on multiple interactions with a single system and observing changes over time, we instead have users engage with multiple different systems in one sitting. The introduction of multiple systems allows users to develop separate mental models and perceptions for each system. The systems vary in context as required by the experimental settings and the contrast between them allows us to understand how contexts impact reliance at the system-level.
@@
Humans perceptions of expressions of (un)certainty are internally consistent Druzdzel (1989) but can vary greatly across subjects Budescu et al. (1988); Chesley (1986); Doupnik and Richter (2003). The variance between subjects is a confounding variable that can be expensive to solve (i.e., recruit thousands of participants). To mitigate this, we have the same participant rely on two different systems, allowing us to robustly measure intra-subject reliance and use that as a relative score of how contextual interactions influence reliance.
@@
Rel-A.I. consists of three core components, a self-incentivized task, epistemic markers from publicly deployed LLMs, and meta-level debrief questions.
@@
Lastly, a single participant will partake in multiple rounds of game play with different agents, each with slightly different interactional contexts. The self-incentivized task can be repeated any number of times, enabling researchers the flexibility to test a variety of contexts.
@@
At the center of this methodology is the integration of epistemic markers elicited from publicly deployed LLMs into the task. We display to users what language models might actually generate in real-world scenarios and measure how often humans will choose to rely on these generations. In our evaluations, the agent always begins their response with an expression of (un)certainty which cues the participants to the LLM’s degree of confidence. These expressions are the most frequently generated expressions as elicited by Zhou et al. (2024) from nine publicly deployed models (text-davinci-003, GPT-3.5-Turbo, GPT-4, LLaMA-2 7B, LLaMA-2 13B, LLaMA-2 70B, Claude-1, Claude-2, Claude- Instant-1), see Table 6 for details. The expressions are then classified into three categories: weakeners—expressions of uncertainty (e.g., ‘‘I think it might be …”, “It could be…”), strengtheners—expressions of certainty (e.g., “I’m certain it’s…”, “Undoubtedly it’s…”), and weakened strengtheners—expressions of moderate certainty (e.g., ”I believe it’s…”, ”Likely it’s…”). [^6] The evaluation of a specific model would be feasible by using the most frequently generated expressions of uncertainty from the model of choice when using the Rel-A.I. approach.
@@ 3.3 Experimental Set Up
With the Rel-A.I. approach, we conduct three experiments, perturbing different interaction variables for each. The first changes the presentation of the mode (RQ1), the second varies the average confidence of the models (RQ2), and the last diversifies the domain of interaction (RQ3).
@@
Each experiment includes 60 or 90 interactions with different models and variations per system are listed in Table 1. Systems and templates are both presented in random order to the human annotators. We launched the task using Prolific and Qualtrics and recruited 50 new participants for each of our three experimental settings. Our task is compliant with internal review board protocols (Appendix A.1). See A.4 for task screenshots and consent form.
Figure 3: Users are based on their change in perception between $A_{control}$ and $A_{exper}$ and observe the reliance rates in each cluster. We see that changes to perceptions of competence, warmth, and humanlikeness are strongly correlated with changes to reliance rates.
4 Varying The Presentation of the Model (RQ1)
The first interactional context we consider is the presentation and anthropomorphism of large language technologies Abercrombie et al. (2023b); Araujo (2018); Colombatto and Fleming (2024). System attributes like names Wagner et al. (2019) as well as human-like language such as greetings and expressions of warmth Bai et al. (2022) are coveted features as they enable more engaging and enriching human-model interaction. However, these changes in presentation hold the potential for encouraging users to unquestioningly rely on system outputs that are prone to errors and hallucinations Cheng et al. (2024); Abercrombie et al. (2023a). As publicly deployed models have been tuned to incorporate helpfulness Bai et al. (2022), we are interested in the impact of greetings on human reliance of epistemic markers.
Experimental Set-Up
Users interact with a control agent and an experimental one ($A_{control}$ and $A_{greet}$). The two agents are identical in response and use of epistemic markers however $A_{greet}$ randomly begin half of its generations with a greeting (see table 2 for list of greetings used). Here, we are interested in how humans behaviors change with the introduction of a simple greeting. Since users have different baseline reliance rates and perceptions, we use $A_{control}$ to normalize values from $A_{greet}$ (i.e., taking the difference between the two).
$\Delta$ Warmth $\Delta$ Competence $\Delta$ Humanlikeness $\Delta$ Reliance
Average Change (%) $\uparrow$33.21* $\uparrow$8.13* $\uparrow$4.61* $\uparrow$1.68*
“I’m happy to help!” $\uparrow$49.17* $\uparrow$7.27* $\uparrow$15.33* $\uparrow$2.87*
“Thank you for the question!” $\uparrow$33.83* $\uparrow$10.43* $\uparrow$6.03* $\uparrow$1.13*
“I’m glad you’re interested in geography!” $\uparrow$44.83* $\uparrow$15.30* $\uparrow$4.37* $\uparrow$2.53*
“I’m here to help.” $\uparrow$25.53* $\uparrow$2.80* $\uparrow$5.30* $\uparrow$0.13
“I can assist with you that.” $\uparrow$12.67* $\uparrow$4.83* $\downarrow$-8.00* $\uparrow$1.73*
Table 2: Change in reliance and perceived characteristics of agents based on the introduction of greetings.
Reliance As Correlated with Perceived Attributes
We cluster users based on their perception of $A_{greet}$’s attributes and observe the reliance rates in each cluster. Figure 3 illustrates that perceptions of warmth, competence, and humanlikeness are all strongly correlated with changes to reliance rates (Pearson’s $r$: 0.78, 0.96, and 0.93 respectively). Notably, there exists a 30% difference in reliance rates based on varying perceptions of competence. Although all three variables explain variance in reliance, an ordinary least regression analysis with all three factors shows that perceived competence is the dominating factor. See A.5 for details.
We also examine the effect of each expression individually (the difference between $A_{control}$ and $A_{exper}$) and observe significant increases in reliance and perceptions of competence, warmth, and humanlikeness (up to 49% increase, Table 2).
Discussion
Work from Cheng et al. (2022) illustrates that the anthropomorphism of chat models in consumer settings leads to increased perceptions of trust, Kim et al. (2024) shows in a medical setting that hedges with personal pronouns are less likely to be relied upon than those without, meanwhile Inie et al. (2024) illustrates the interaction effects between anthropomorphism and product type. Our work adds to prior work and the simple introduction of greetings can have a consequential influence on perceptions of model warmth, competence, and humanlikeness, and this in turn is correlated with human reliance behaviors. Together, it suggests that a more careful use of these cues may be warranted.
Lastly, in human perceptions of humans, warmth and competence have a hydraulic relationship where when perceptions of one go up, the other goes down Cuddy et al. (2011). However, our results show that as perceived warmth increases, perceived competence does not decrease. In fact, we see a slight increase. These findings potentially highlight that the hydraulic relationship of warmth and competence may apply differently to perceptions of AI agents.
5 Influence of Prior Interactions (RQ2)